a garden is shaped like a rectangle whose perimeter is 66 ft. The length is 9 ft more than the width. Find the length and the width.
Answer by JulietG(1812)

A perimeter is a fence around. 2 lengths and 2 width = perimeter (2W + 2L = P)
L = W + 9
Substitute the value of L from the second equation into the first.
2W + 2(W+9) = 66
Distribute
2W + 2W + 18 = 66
Add
4W + 18 = 66
Subtract 18 from each side
4W = 48
Divide each side by 4
W = 12
.
If the width is 12, then the length is 12+9, or 21.